OpenStudy (bobobox):

I am completely lost, I require help solve for x: −3|2x + 6| = −12 (1 point) (a) x = 1 and x = 5 (b) x = −1 and x = −5 (c) x = −9 and x = 3 (d) No solution

OpenStudy (bobobox):

@satellite73

OpenStudy (bobobox):

@KyleWise16

OpenStudy (tkhunny):

Why are you lost? Use all the stuff you already know. −3|2x + 6| = −12 Divide by -3 |2x + 6| = 4 Think for a moment. Absolute values are not ever negative. This one is good. If it said, |2x+6| = -4, you could just quit. There is no solution. Continue. Factor out the common 2. 2|x + 3| = 4 Divide by 2 |x + 3| = 2 Okay, this is where it MIGHT be a little different. Everything up until now, you should have been able to do. Parentheses or absolute values, it is the same up to this point. Now what? Can you think of any numbers whose absolute value is 2?

OpenStudy (bobobox):

I thought absolute value is how far away from zero the number is so I think either 2 or -2

OpenStudy (tkhunny):

Perfect. Solve!! x+3 = 2 or x+3 = -2 Go!

OpenStudy (bobobox):

x+3=-2 i would think because two couldn't be made by adding three and something else so x must be a negitive number

OpenStudy (tkhunny):

?? It's both of them. There are two solutions.